# Positive concave deep equilibrium models

### v1

@@ 5 Experiments
The experiments were carried out on three commonly known computer vision datasets: MNIST, SVHN, and CIFAR-10. The proposed pcDEQ models were compared with competitive approaches for which the uniqueness of the fixed point is mathematically established, namely: monotone operator deep equilibrium models (monDEQ), neural ordinary differential equations (NODE), and augmented neural differential equations (ANODE). The comparison was performed according to the results reported in (Dupont et al., 2019; Winston & Kolter, 2020). For pcDEQ models, the fixed point $z^{\star}$ is computed using the standard fixed point iteration in (2). The stopping criterion of fixed point iteration is based on the relative error, calculated as $\frac{||z_{k+1}-z_{k}||}{||z_{k+1}||}\leq\epsilon$, where |||||$\cdot$||| | ⋅ | | is a Frobenius norm and $\epsilon$ is a tolerance. In our experiments, the tolerance $\epsilon$ was set to $41e-$ 4.
@@ 5.1 Architecture setup
In experiments, we consider three types of pcDEQ models with four activation functions from Remark 4.1. The first type of network is based on using one linear pcDEQ layer. The second type of network is based on using one convolutional pcDEQ layer, and the last type is based on the use of three convolutional pcDEQ layers, between which the explicit downsampling layers occur. The architectural details are discussed in Appendix B, and the experimental hyperparameters for each network in Appendix C. Networks using a single linear pcDEQ layer have “L” suffix in their name. A similar scenario is for networks with a single convolutional pcDEQ layer, to network name suffix “SC” is added. For networks with three convolutional pcDEQ layers, the suffix “MC” is added to the network name.
@@ 5.2 Results
Table 1: Test accuracies of pcDEQ models on MNIST dataset compared with NODE, ANODE and monDEQ.
@@
monDEQ-MC 81K 99.0
pcDEQ-1-L-ReLU6 70K 98.2
@@
pcDEQ-1-SC-Tanh 69K 99.1
pcDEQ-1-SC-Softsign 69K 99.0
pcDEQ-2-SC-Sigmoid 69K 99.0
pcDEQ-1-MC-ReLU6 41K 99.4
@@
pcDEQ-1-MC-Softsign 41K 99.3
pcDEQ-2-MC-Sigmoid 41K 98.6
Table 2: Test accuracies of pcDEQ models on SVHN dataset compared with NODE, ANODE and monDEQ.
@@
monDEQ-MC 170K 92.4
pcDEQ-1-SC-ReLU6 165K 87.9
pcDEQ-1-SC-Tanh 165K 88.3
pcDEQ-1-SC-Softsign 165K 88.5
pcDEQ-2-SC-Sigmoid 165K 87.6
pcDEQ-1-MC-ReLU6 131K 93.1
pcDEQ-1-MC-Tanh 131K 91.9
pcDEQ-1-MC-Softsign 131K 92.7
pcDEQ-2-MC-Sigmoid 131K 90.7
Table 3: Test accuracies of pcDEQ models on CIFAR-10 dataset compared with NODE, ANODE and monDEQ; * with data augmentation.
@@
monDEQ-MC 170K 72.0
monDEQ-MC* 1M 89.0
pcDEQ-1-SC-ReLU6 165K 76.4
pcDEQ-1-SC-Tanh 165K 77.1
pcDEQ-1-SC-Softsign 165K 76.6
@@
pcDEQ-1-MC-ReLU6 131K 78.3
pcDEQ-1-MC-ReLU6* 661K 89.1
pcDEQ-1-MC-Tanh 131K 77.2
pcDEQ-1-MC-Tanh* 661K 88.3
pcDEQ-1-MC-Softsign 131K 76.6
pcDEQ-1-MC-Softsign* 661K 89.3
pcDEQ-2-MC-Sigmoid 131K 71.4
pcDEQ-2-MC-Sigmoid* 661K 85.8
@@
Figure 2: Test accuracies during training for the pcDEQ model with a single convolutional layer.
@@ 5.3 Convergence analysis
Figure 3: Number of fixed point iterations for computing the fixed point in forward and backward passes for the pcDEQ model with a single convolutional layer.
@@ 5.4 Lipschitz continuity
In this section, we show empirically that the assumptions used to determine the uniqueness of the fixed point are weaker compared to the standard assumptions in convex analysis. In general, to determine the uniqueness of the fixed point based on the Banach fixed point theorem, we need information about a Lipschitz constant of the function. In the case of neural networks, if the used activation functions are nonexpansive (such as those in Remark 4.1), then a Lipschitz constant $L$ is upper bounded by the product of spectral norms of weight operators (Combettes & Pesquet, 2020) as follows:
@@
where $m$ is the number of weight operators and |||||$\cdot$||| | ⋅ | | is spectral norm.
@@
Figure 4: Largest singular value of pcDEQ linear layer.
@@ 6 Limitations
We admit that the proposed approach is quite restrictive due to the non-negative restriction of weights and concave activation functions. However, in order to have some theoretical guarantees, such as the uniqueness of the fixed point and the guarantees of finding it by a computationally efficient method, restrictions are necessary. Now it is possible to discuss which restrictions on layers better generalize in practice. One can think that positive concavity is very restrictive, but the strong monotonic assumption in monDEQ is also very restrictive. The nonnegative restrictions on weights are much simpler to implement than constraining weights to be strongly monotone, as it is the case in monDEQ models. The family of mappings with similar properties to SI mappings can be much larger. There exists some work on the extension of the classical linear Perron-Frobenius theory to matrices with some negative entries (Noutsos, 2006). Extending these results to nonlinear settings is a possible venue for improving pcDEQ models.
@@ 7 Conclusions
This study provides the foundation for extending the proposed method to a larger and less restricted class of mappings with guarantees similar to SI mappings. As an example, one may aim to weaken further the assumptions used to construct pcDEQ models to obtain more versatile models, capable of incorporating a wider class of weight operators and activation functions in DEQ layers. Such a direction seems feasible in view of the recent results on the existence and shape of the fixed point sets of (subhomogeneous) weakly scalable (WSI) neural networks provided in (Piotrowski et al., 2021). Another promising future research in the direction proposed in this paper should focus on providing even stronger guarantees of convergence rate compared with the geometric convergence of the fixed point iteration of pcDEQ models established in this paper. Indeed, the empirical convergence analysis provided in Section 5.3 suggests that the actual rate of convergence may actually be linear and, for certain pcDEQ architectures, even quadratic.
@@ Appendix A Known results
(Ortega & Rheinboldt, 2000)[Chapter 9] Let $(x_{k}:k\in\mathbb{N})$ be the sequence in $\mathbb{R}^{n}$, then $(x_{k}:k\in\mathbb{N})$ converges geometrically to $x^{\star}\in\mathbb{R}^{n}$ with a rate $c\in[0,1)$ and a constant $\gamma>0$ if
@@
(Cavalcante et al., 2019) Let $f\colon\mathbb{R}^{n}_{+}\to\text{int}(\mathbb{R}^{n}_{+})$ be a SI mapping. The asymptotic mapping associated with $f$ is the mapping defined by
$f_{\infty}:\mathbb{R}^{n}_{+}\to\mathbb{R}^{n}_{+}:z\mapsto\lim_{p\to\infty}\frac{1}{p}f(px).$ (11)
We recall that the above limit always exists and that the resulting asymptotic mapping $f_{\infty}$ is positively homogeneous in the sense that $(\forall\alpha>0)$ $(\forall x\in\mathbb{R}^{n}_{+})$ $f_{\infty}(\alpha x)=\alpha f_{\infty}(x).$ See (Cavalcante et al., 2019) for proofs of these facts.
@@
The (nonlinear) spectral radius of an SI mapping is defined as the largest eigenvalue of the corresponding asymptotic mapping (Cavalcante et al., 2019; Oshime, 1992) given by
$\begin{split}p(f_{\infty}):=\text{max}\{\lambda\in\mathbb{R}_{+}\;|\;\exists x\in\mathbb{R}^{n}_{+}\setminus\{0\}\;s.t.\;f_{\infty}(x)=\lambda x\}\in\mathbb{R}_{+}.\end{split}$ (12)
@@
[…]
@@
(Cavalcante et al., 2015) If $f\colon\mathbb{R}^{n}_{+}\to\mathbb{R}^{n}_{+}$ is an NC mapping or $f\colon\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ is a PC mapping in the sense of Definition 3.4, then $f$ is monotonic.
@@
(Cavalcante et al., 2015, 2019) If $f:\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ is a PC mapping, then $f$ is an SI mapping.
@@ B.2 Architecture with single pcDEQ convolutional layer
B.3 Architecture with multiple pcDEQ convolutional layer
@@ Appendix C Experimental details and hyperparameters
As mentioned in the paper, to compute fixed point in pcDEQ layers, the standard fixed point iteration was used. We especially do not used any more sophisticated iterative methods such as Anderson acceleration or Broyden’s method, because for such methods the guarantees of convergence to fixed point for PC mappings have not been proved in the literature. For all networks, the AdamW (Loshchilov & Hutter, 2017) optimizer was used with a batch size of 64. All other hyperparameters for each dataset and architecture are shown in Tables 4, 5 and 6.
@@
pcDEQ-1-L-ReLU6 80 40 0.0007 30 0.1 0.02
pcDEQ-1-L-Tanh 80 40 0.0007 30 0.1 0.02
pcDEQ-1-L-Softsign 80 40 0.0007 30 0.1 0.02
pcDEQ-2-L-Sigmoid 80 40 0.0007 30 0.1 0.02
@@ D.1 MNIST
Figure 8: Test accuracies, number of fixed point iterations in forward and backward passes during training for pcDEQ models with single linear pcDEQ layer and three convolutional pcDEQ layers on MNIST dataset.
@@ D.2 SVHN
Figure 9: Test accuracies, number of fixed point iterations in forward and backward passes during training for pcDEQ models with multiple convolutional pcDEQ layers on SVHN dataset.
@@ D.3 CIFAR-10
Figure 10: Test accuracies, number of fixed point iterations in forward and backward passes during training for pcDEQ models with three convolutional pcDEQ layers with and without data augmentation on CIFAR-10 dataset.

### v2

@@ 5 Experiments
The experiments were carried out on three commonly known computer vision datasets: MNIST, SVHN, and CIFAR-10. The proposed pcDEQ [^1] The Pytorch source code of pcDEQs along with examples of use is available at the following link: https://github.com/mateuszgabor/pcdeq models were compared with competitive approaches for which the uniqueness of the fixed point is mathematically established, namely: monotone operator deep equilibrium models (monDEQ) (Winston & Kolter, 2020), neural ordinary differential equations (NODE) (Chen et al., 2018), and augmented neural differential equations (ANODE) (Dupont et al., 2019). The comparison was performed according to the results reported in (Dupont et al., 2019; Winston & Kolter, 2020). For pcDEQ models, the fixed point $z^{\star}$ is computed using the standard fixed point iteration in (2). The stopping criterion of fixed point iteration is based on the relative error, calculated as $\frac{||z_{k+1}-z_{k}||}{||z_{k+1}||}\leq\epsilon$, where |||||$\cdot$||| | ⋅ | | is a Frobenius norm and $\epsilon$ is a tolerance. In our experiments, the tolerance $\epsilon$ was set to $41e-$ 4.
@@ 5.1 Architecture Setup
In experiments, we consider three types of pcDEQ models with four activation functions from Remark 4.1. The first type of network is based on using one linear pcDEQ layer. The second type of network is based on using one convolutional pcDEQ layer, and the last type is based on the use of three convolutional pcDEQ layers, between which the explicit downsampling layers occur. The architectural details are discussed in Appendix B, and the experimental hyperparameters for each network in Appendix C. Networks using a single linear pcDEQ layer have “L” suffix in their name. A similar scenario is for networks with a single convolutional pcDEQ layer, to network name suffix “SC” is added. For networks with three convolutional pcDEQ layers, the suffix “MC” is added to the network name. The suffix “MT” in monDEQs refers to the multi-tier architecture used in (Winston & Kolter, 2020).
@@ 5.2 Results
Table 1: Test accuracies of pcDEQ models averaged over five runs on MNIST dataset compared with NODE, ANODE and monDEQ; $\dagger$ as reported in (Dupont et al., 2019); $\ddagger$ as reported in (Winston & Kolter, 2020).
@@
monDEQ-MT 81K 99.0
pcDEQ-1-L-ReLU6 70K 98.1
@@
pcDEQ-1-SC-Tanh 69K 99.2
pcDEQ-1-SC-Softsign 69K 99.1
pcDEQ-2-SC-Sigmoid 69K 98.9
pcDEQ-1-MC-ReLU6 41K 99.3
@@
pcDEQ-1-MC-Softsign 41K 99.2
pcDEQ-2-MC-Sigmoid 41K 98.7
Table 2: Test accuracies of pcDEQ models averaged over five runs on SVHN dataset compared with NODE, ANODE and monDEQ; $\dagger$ as reported in (Dupont et al., 2019); $\ddagger$ as reported in (Winston & Kolter, 2020).
@@
monDEQ-MT 170K 92.4
pcDEQ-1-SC-ReLU6 165K 88.0
pcDEQ-1-SC-Tanh 165K 88.1
pcDEQ-1-SC-Softsign 165K 88.4
pcDEQ-2-SC-Sigmoid 165K 87.3
pcDEQ-1-MC-ReLU6 131K 93.0
pcDEQ-1-MC-Tanh 131K 92.3
pcDEQ-1-MC-Softsign 131K 92.3
pcDEQ-2-MC-Sigmoid 131K 91.5
Table 3: Test accuracies of pcDEQ models averaged over five runs on CIFAR-10 dataset compared with NODE, ANODE and monDEQ; $\dagger$ as reported in (Dupont et al., 2019); $\ddagger$ as reported in (Winston & Kolter, 2020); * with data augmentation.
@@
monDEQ-MT 170K 72.0
monDEQ-MT* 1M 89.0
pcDEQ-1-SC-ReLU6 165K 76.3
pcDEQ-1-SC-Tanh 165K 76.6
pcDEQ-1-SC-Softsign 165K 76.4
@@
pcDEQ-1-MC-ReLU6 131K 78.2
pcDEQ-1-MC-ReLU6* 661K 89.2
pcDEQ-1-MC-Tanh 131K 76.5
pcDEQ-1-MC-Tanh* 661K 88.5
pcDEQ-1-MC-Softsign 131K 77.1
pcDEQ-1-MC-Softsign* 661K 89.0
pcDEQ-2-MC-Sigmoid 131K 71.0
pcDEQ-2-MC-Sigmoid* 661K 85.6
@@
Figure 2: Test accuracies during training for the pcDEQ model with a single convolutional layer over five experiment runs.
@@ 5.3 Convergence Analysis
Figure 3: Number of fixed point iterations for computing the fixed point in forward and backward passes for the pcDEQ model with a single convolutional layer over five experiment runs.
@@ 5.4 Lipschitz Continuity
In this section, by considering the standard Euclidean metric space, we show empirically that the assumptions used to determine the uniqueness of the fixed point are weaker compared to the standard assumptions in convex analysis. In general, to determine the uniqueness of the fixed point based on the Banach fixed point theorem, we need information about a Lipschitz constant of the function, and, in the machine learning literature, we often consider the standard Euclidean space. In the case of neural networks, if the used activation functions are nonexpansive (such as those in Remark 4.1), then a Lipschitz constant $L$ w.r.t. the standard Euclidean space is upper bounded by the product of spectral norms of weight operators (Combettes & Pesquet, 2020) as follows:
@@
where $m$ is the number of weight operators and |||||$\cdot$||| | ⋅ | | is the spectral norm.
@@
Figure 4: Largest singular value of pcDEQ linear layer over five experiment runs.
@@ 6 Limitations
The proposed approach may appear restrictive due to the non-negativity constraint on weights and the use of concave activation functions. However, these restrictions are essential to ensure theoretical guarantees, such as the uniqueness of the fixed point and the convergence of the standard fixed point iteration. It is important to note that existing approaches offering similar guarantees, such as monDEQ models, also impose constraints to ensure that the mappings are strongly monotone in the sense used in convex analysis. The advantage of our approach is that the imposed restrictions are simpler to implement in practice compared to those based on monotone operator theory.
@@ 7 Conclusions
This study provides the foundation for extending the proposed method to a larger and less restricted class of mappings with guarantees similar to SI mappings. As an example, one may aim to weaken further the assumptions used to construct pcDEQ models to obtain more versatile models, capable of incorporating a wider class of weight operators and activation functions in DEQ layers. Such a direction seems feasible in view of the recent results on the existence and shape of the fixed point sets of (subhomogeneous) weakly standard interference (WSI) neural networks provided in (Piotrowski et al., 2024). Another promising future research in the direction proposed in this paper should focus on providing even stronger guarantees of convergence rate compared with the geometric convergence of the fixed point iteration of pcDEQ models established in this paper. Indeed, the empirical convergence analysis provided in Section 5.3 suggests that the actual rate of convergence may actually be linear.
Acknowledgments
This work is supported by the National Center of Science under Preludium 22 project No. UMO-2023/49/N/ST6/02697 (Mateusz Gabor), National Centre for Research and Development of Poland (NCBR) under grant EIG CONCERT-JAPAN/04/2021, and by the Federal Ministry of Education and Research of Germany under grant 01DR21009 and the programme “Souverän. Digital. Vernetzt.” Joint project 6G-RIC, project identification numbers: 16KISK020K and 16KISK030.
@@ Appendix A Known Results
(Ortega & Rheinboldt, 2000)[Chapter 9] Let $(x_{k}:k\in\mathbb{N})$ be a sequence in $\mathbb{R}^{n}$, then $(x_{k}:k\in\mathbb{N})$ converges geometrically to $x^{\star}\in\mathbb{R}^{n}$ with a rate $c\in[0,1)$ and a constant $\gamma>0$ if
@@
(Cavalcante et al., 2019; Oshime, 1992) Let $f\colon\mathbb{R}^{n}_{+}\to\text{int}(\mathbb{R}^{n}_{+})$ be an SI mapping in the sense of Definition 3.2. The asymptotic mapping associated with $f$ is the mapping defined by
$f_{\infty}:\mathbb{R}^{n}_{+}\to\mathbb{R}^{n}_{+}:z\mapsto\lim_{p\to\infty}\frac{1}{p}f(pz).$ (11)
We recall that the above limit always exists and that the resulting asymptotic mapping $f_{\infty}$ is positively homogeneous; i.e., $(\forall\alpha>0)$ $(\forall z\in\mathbb{R}^{n}_{+})$ $f_{\infty}(\alpha x)=\alpha f_{\infty}(z).$
@@
The (nonlinear) spectral radius of an SI mapping is defined as the largest eigenvalue of the corresponding asymptotic mapping (Cavalcante et al., 2019; Oshime, 1992), and it is given by
$\begin{split}p(f_{\infty}):=\text{max}\{\lambda\in\mathbb{R}_{+}\;|\;\exists z\in\mathbb{R}^{n}_{+}\setminus\{0\}\;s.t.\;f_{\infty}(z)=\lambda z\}\in\mathbb{R}_{+}.\end{split}$ (12)
@@
(Cavalcante et al., 2019) Let $f\colon\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ be an SI mapping. Then $\text{Fix}(f)\neq\emptyset$ if and only if $p(f_{\infty})<1$. Furthermore, if a fixed point exists, then it is positive and unique.
@@
(Cavalcante et al., 2016) If $f\colon\mathbb{R}^{n}_{+}\to\mathbb{R}^{n}_{+}$ is an NC mapping or $f\colon\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ is a PC mapping in the sense of Definition 3.4, then $f$ is monotonic.
@@
(Cavalcante et al., 2016, 2019) If $f:\mathbb{R}^{n}_{+}\to\textup{int}(\mathbb{R}^{n}_{+})$ is a PC mapping, then $f$ is an SI mapping.
@@ B.2 Architecture with Single pcDEQ Convolutional Layer
B.3 Architecture with Multiple pcDEQ Convolutional Layers
@@ Appendix C Experimental Details and Hyperparameters
In our experiments, we used three commonly known computer vision datasets: MNIST, SVHN, and CIFAR-10. MNIST dataset consists 70,000 grayscale handwritten digit images. SVHN dataset consists of 99,289 RGB digit images from house numbers. CIFAR-10 consists of 60,000 RGB images of 10 classes. For the MNIST dataset, the images have dimensions of $28\times 28$ pixels, and for the SVHN and CIFAR-10 $32\times 32$ pixels. The statistics of the datasets are shown in Table 4.
Table 4: Dataset statistics.
Dataset #Train examples #Test examples
MNIST 60,000 10,000
SVHN 73,257 26,032
CIFAR-10 50,000 10,000
As mentioned in the paper, to compute fixed point in pcDEQ layers, the standard fixed point iteration was used. We especially do not use any more sophisticated iterative methods such as Anderson acceleration or Broyden’s method, because for such methods the guarantees of convergence to fixed point for PC mappings have not been proved in the literature. For all networks, the AdamW (Loshchilov & Hutter, 2017) optimizer was used with a batch size of 64. All other hyperparameters for each dataset and architecture are shown in Tables 5, 6 and 7. The expeeriments with pcDEQs were run five times.
@@
pcDEQ-1-L-ReLU6 80 40 0.001 30 0.1 0.02
pcDEQ-1-L-Tanh 80 40 0.001 30 0.1 0.02
pcDEQ-1-L-Softsign 80 40 0.001 30 0.1 0.02
pcDEQ-2-L-Sigmoid 80 40 0.001 30 0.1 0.02
@@ D.1 MNIST
Figure 8: Test accuracies, number of fixed point iterations in forward and backward passes during training for pcDEQ models with single linear pcDEQ layer and three convolutional pcDEQ layers (average of forward and backward passes of three pcDEQ layers) on MNIST dataset over five experiment runs.
@@ D.2 SVHN
Figure 9: Test accuracies, number of fixed point iterations in forward and backward passes (average of three pcDEQ layers) during training for pcDEQ models with multiple convolutional pcDEQ layers on SVHN dataset over five experiment runs.
@@ D.3 CIFAR-10
Figure 10: Test accuracies, number of fixed point iterations in forward and backward passes (average of three pcDEQ layers) during training for pcDEQ models with three convolutional pcDEQ layers with and without data augmentation on CIFAR-10 dataset over five experiment runs.